A station emits wave of wavelength, Lamda at P. If the planck's constant is (h=6.63×10^-34 js) and the speed of light is (C=3×10^8 ms-1), what's the rate of emission of photon
1
Expert's answer
2019-06-12T11:16:52-0400

the rate of emission of photon equals to

"R = \frac{P}{\epsilon} = \frac{P\lambda}{hc} = 5\cdot 10^{24}P\lambda"

photon/s.